# Efficient 3D Bayesian Full Waveform Inversion and Analysis of Prior Hypotheses

### v1

@@ 6 ACKNOWLEDGMENTS
We thank the Edinburgh Imaging Project (EIP - https://blogs.ed.ac.uk/imaging/) sponsors (BP and TotalEnergies) for supporting this research.
Appendix A Inversion results using a Uniform Prior Distribution
As displayed in Figure 4a, the true velocity model used in the main text is a part of the 3D overthrust model [1], which contains $63\times 101\times 101$ grid cells in Z (depth), Y and X directions, respectively, with a cell size of 50 m in each direction. We deploy 81 sources (red stars) and 10,201 receivers (white dots) at the surface with spacings of 500 m and 50 m, respectively. Data are generated by solving the 3D acoustic wave equation using a time-domain pseudo-analytical method [4] with an Ormsby wavelet for 6 s. We define a Uniform prior distribution for the velocity values at different depths, with upper and lower bounds shown in Figure 4b.
Figures 5 and 6, respectively, display the inversion results obtained using low and intermediate frequency waveform data, performed using physically structured variational inference [PSVI – 14]. Similarly to the inversion results presented in Figure 1 in the main text, the top two rows display two vertical sections at locations Y = 2.5 km and X = 2.5 km, and bottom row illustrates one horizontal section at Z = 1.25 km. From left to right, each column shows three true velocity sections (marked by black dashed lines in Figure 1a), the corresponding inverted average velocity, the standard deviation, and the relative error maps of the posterior distribution, where the relative error is the absolute difference between the true and mean values divided by the standard deviation at each point. Comparing these results with those using the high frequency data we observe that as the frequency of waveform data used for inversion increases, the spatial resolution becomes higher, and velocity estimates become more accurate. The overall relative errors are within 3 standard deviations, which might be expected of the true posterior probability distribution.
Figure 4: (a) True velocity model. Source and receiver locations are indicated by red stars and white dots at the surface (the latter are so close together that they may be observed as a pale haze). Dashed black lines mark the locations of three 2D slices discussed in the text. (b) Upper and lower bounds of a Uniform prior distribution at each depth.
Figure 5: Low frequency 3D variational Bayesian FWI results obtained using PSVI and the Uniform prior distribution. Top two rows show two vertical sections at Y = 2.5km and X = 2.5km, and bottom row shows a horizontal section at Z = 1.25km. Panels (a) – (d) display the true velocity model, the mean, the standard deviation and the relative error maps of the posterior distribution.
@@ Appendix B Verification of Variational Prior Replacement
In this section, we test the effectiveness and accuracy of variational prior replacement [VPR – 12] for 3D Bayesian FWI by comparing the posterior solutions obtained using VPR to those obtained from independent Bayesian inversions. For VPR, we consider the Uniform prior distribution defined in the main text as the old prior pdf, which is then removed from the old posterior distribution (displayed in Figures 1a to 1d in the main text) and replaced by a smoothed prior distribution with $\sigma_{s}=1000$ (see details in the main text). The corresponding results are displayed in Figure 7. We then perform an independent variational Bayesian inversion using the same smoothed prior distribution. Similarly to the Uniform prior case, the inversion is performed using the same three frequency bands of waveform data with the same number of (5200) forward simulations. Figure 8 illustrates the results. The main features of these two results are roughly consistent, indicating that VPR can replace the old with the new prior information in the inversion results without solving a Bayesian inverse problem repeatedly from scratch. Some small discrepancies between these two sets of results remain, especially in the deeper parts of the model. This is partly because VPR introduces a variational distribution $q_{new}(\mathbf{m})$ to approximate the new posterior pdf $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$, rather than calculating the true new posterior distribution, as expressed in the fifth line in equation 4 in the main text [8, 12]. In addition, this also indicates that VPR might not be as accurate for this extremely high dimensional and complex 3D Bayesian inverse problem, compared to previous tests in 2D [12]. Nevertheless, the result in Figure 7 is obtained almost for free, compared to the huge computational cost typically involved in 3D forward simulation [15] and Bayesian inversion [11]. In the main text, we introduce a fining tune scheme to improve the inversion results displayed in Figure 7.
@@
Appendix C Posterior Marginal pdf’s using Smoothed Priors
Figure 9 compares posterior marginal pdf’s of two vertical velocity profiles at locations X = 1km (top row of Figure 9) and X = 3km (bottom row), marked by red dashed lines in Figure 7a. Figure 9a displays the posterior marginal pdf’s obtained from the Uniform prior distribution, and Figures 9b – 9h show those from the smoothed prior pdf’s – the smoothness values $\sigma_{s}$ are noted in the title. In each figure, the red line shows the true velocity profile and black line displays the inverted mean velocity profile. We observe that with the increase of the magnitude of smoothness injected by the prior information, the posterior marginal pdf’s become smoother with fewer spatial variations (thus lower uncertainties), similarly to the conclusions drawn in the main text.
Figure 9: Posterior marginal pdf’s at two vertical locations marked by red dashed lines in Figure 7a. (a) shows results obtained using the Uniform prior pdf, and (b) – (h) correspond to those using the smoothed prior pdf’s with different $\sigma_{s}$ values.
Appendix D Inversion results using Geological Prior information
In addition to the smoothed prior distributions mentioned in the main text, we also consider another set of prior distributions with real-geology informed inter-parameter correlation information. This is achieved by selecting a set of realistic geological images and then subsampling them using a smaller window with a predefined window size. These subimages are used to calculate a local correlation matrix to define prior correlations between pairs of model parameters. More details about this geological prior distribution can be found in [12]. In this section, we consider 5 such geological prior distributions with different local window sizes of $2\times 2\times 2$, $4\times 4\times 4$, $6\times 6\times 6$, $8\times 8\times 8$ and $10\times 10\times 10$, respectively. Again, the corresponding posterior distributions are calculated using VPR by replacing the old Uniform prior distribution by each defined geological prior distribution. Figure 10 displays statistics of the corresponding posterior pdf’s, where the sizes of the correlation windows used are denoted on the left side of Figure 10. Figure 10a represents the results obtained using the Uniform prior distribution. We observe that larger prior correlation windows generally result in better inversion results that are more similar to the true velocity model, since a larger window can define better prior correlation information, which improves the inversion results. However, a larger window also has higher memory requirements. These inversion results are then used to construct a Bayesian L-curve, as displayed by the dashed blue line in Figure 3b in the main text. From left to right along this dashed blue line, the correlation window size decreases from $10\times 10\times 10$ (top-left point) to $0\times 0\times 0$ (bottom-right point), the latter being the Uniform prior pdf with no prior correlation.

### v2

@@ 6 ACKNOWLEDGMENTS
We thank the Edinburgh Imaging Project (EIP - https://blogs.ed.ac.uk/imaging/) sponsors (BP and TotalEnergies) for supporting this research. We also thank Dr. Xin Zhang from China University of Geosciences in Beijing for providing insightful discussion about three variational inversion methods (ADVI, SVGD and sSVGD) and the corresponding inversion results.
Appendix A Inversion results using low and intermediate frequency waveform data
Figures 10 and 11, respectively, display the inversion results obtained using low and intermediate frequency waveform data, performed using physically structured variational inference [PSVI – 52]. Similarly to the inversion results displayed in Figure 4 in the main text, the top two rows display two vertical sections at locations Y = 2.5 km and X = 2.5 km, and the bottom row illustrates one horizontal section at Z = 1.25 km. From left to right, each column shows three true velocity sections (marked by black dashed lines in Figure 1a), the corresponding inverted average velocity, the standard deviation, and the relative error maps of the posterior distribution, where the relative error is the absolute difference between the true and mean values divided by the standard deviation at each point. Comparing these results with those using the high frequency data (Figure 4) we observe that as the frequency of waveform data used for inversion increases, the spatial resolution becomes higher, and velocity estimates become more accurate. The overall relative errors are within 3 standard deviations, which might be expected of the true posterior probability distribution.
Figure 10: Low frequency 3D variational Bayesian FWI results obtained using PSVI and a uniform prior distribution. Top two rows show two vertical sections at Y = 2.5km and X = 2.5km, and bottom row shows a horizontal section at Z = 1.25km. Panels (a) – (d) display the true velocity model, the mean, the standard deviation and the relative error maps of the posterior distribution.
@@ Appendix B Verification of Variational Prior Replacement
In this Appendix, we test the effectiveness and accuracy of variational prior replacement [VPR – 53] for 3D Bayesian FWI by comparing the posterior solutions obtained using VPR to those obtained from independent Bayesian inversions. For VPR, we consider the uniform prior distribution defined in the main text as the old prior pdf, which is then removed from the old posterior distribution (displayed in Figure 4 in the main text) and replaced by a smoothed prior distribution with $\sigma_{s}=1000$ (see details in the main text). The corresponding results are displayed in Figure 12. We then perform an independent variational Bayesian inversion using the same smoothed prior distribution. Similarly to the uniform prior case, the inversion is performed using the same three frequency bands of waveform data with the same number of forward simulations. Figure 13 illustrates the results. The main features of these two results are roughly consistent, indicating that VPR can replace the old with the new prior information in the inversion results without solving a Bayesian inverse problem repeatedly from scratch. Some small discrepancies between these two sets of results remain, especially in the deeper parts of the model. This is partly because VPR introduces a variational distribution $q_{new}(\mathbf{m})$ to approximate the new posterior pdf $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$, rather than calculating the true new posterior distribution, as expressed in the fifth line in equation 5 in the main text [42, 53]. In addition, this also indicates that VPR might not be as accurate for this extremely high dimensional and complex 3D Bayesian inverse problem, compared to previous tests in 2D [53]. Nevertheless, this result is obtained almost for free, compared to the huge computational cost typically involved in 3D forward simulation [55] and Bayesian inversion [50]. In the main text, we introduce a fine tuning scheme to improve the inversion results displayed in Figure 12.